OpenStudy (anonymous):

how do we solve this "t-3/2 = 75" please i need help!!

OpenStudy (anonymous):

\[t - \frac{3}{2} =75\]STEP I) Add 3/2 to both sides\[t-\frac{3}{2}+\frac{3}{2}=75+\frac{3}{2}\]so you get\[t = 75+\frac{3}{2}\]Simplify to get the answer.